# AesExpert: Towards Multi-modality Foundation Model for Image Aesthetics Perception

### v1

@@ 3.3. Generating Instruction-following Data
Aesthetic Question Answering: In addition to directly adopting the AesFeedback as aesthetic instruction-following data, we further propose a GPT-assist approach to refine the aesthetic critiques and assemble a larger-scale aesthetic question-answering subset (AesVQA). Inspired by COCO-VQA (Lin et al., 2015) and ShareGPT (Chen et al., 2023a), we prompt GPT-4 to generate diverse-style open-ended questions and provide corresponding brief answers, where both the questions and answers are based on human aesthetic descriptions. To ensure that MLLMs can handle diverse queries during the interaction, the questions are generated to cover commonly used question types. Specifically, Yes-or-No-style questions are straightforward queries that demand a simple yes or no as an answer, which are mainly used to provide a clear, binary response. What-style questions are leveraged to measure more comprehensive and complex aesthetic perception (e.g. various aesthetic attributes). How-style questions are used to ask more details about aesthetic emotion, aesthetic attributes, and improvement suggestions; Why-style questions are employed to explore the foundational aspects of what makes images aesthetically appealing or unappealing, revealing the rationale of beauty and visual appeal. To ensure the diversity of instructions, following the existing works (Liu et al., 2023a; Lu et al., 2022), besides the direct answers, we also generate several distracting answers for the questions and convert them into an additional multi-choice question format (Wu et al., 2024b). Through the above process, we obtain the final AesMMIT dataset, which consists of 409K multi-typed instructions, aiming to activate stronger aesthetic capabilities. Figure 4 shows the frequently occurring words in the proposed AesMMIT dataset. It can be observed that most nouns, adjectives and adverbs are related to image aesthetic description, which is quite different from common semantic-based tasks.
Figure 4. Statistics of top-20 frequently-used words in AesMMIT database.
@@ 4.1. Model Architecture
The proposed AesExpert model follows the design of LLaVA-1.5 (Liu et al., 2023a), which includes three components: (1) A visual model based on the CLIP-ViT-L14 (Radford et al., 2021) with an input size of $\times×336$, which converts the input image into 576 tokens. (2) A visual-language projector based on two-layer multi-layer perception (MLP), which is employed to connect the visual modality and language modality. (3) A language model based on the open-source Vicuna-v1.5 (Zheng et al., 2024), which is used to interpret instructions and generate responses. In this work, we build the AesExpert models based on two variants of LLaVA-1.5, including LLaVA-1.5-7B and LLaVA-1.5-13B. In addition, we also introduce the mPLUG-Owl2 (Ye et al., 2023b) as the backbone to validate the aesthetic perception performance based on the AesMMIT dataset. The structures for these multi-modality foundation models are summarized in Table 2.
@@
Q-Instruct 66.03% 74.48% 73.68% 68.09% 72.61% 68.64% 83.86% 75.86% 80.00% 72.68% 52.86% 1.020 75.04% 8
GLM 55.77% 54.61% 51.25% 48.94% 52.96% 53.13% 70.97% 44.83% 55.29% 53.96% 37.79% 0.932 59.48% 9
MiniGPT-4 39.42% 41.31% 42.67% 44.33% 41.93% 39.78% 38.71% 24.14% 39.04% 39.35% 38.57% 0.999 54.94% 10
mPLUG-Owl2 60.90% 70.57% 68.30% 62.77% 67.89% 65.60% 77.42% 65.52% 78.07% 69.89% 50.57% 1.182 47.38% 11
LLaVA-1.5-7B 53.85% 70.16% 67.40% 59.93% 66.32% 62.49% 80.65% 75.85% 78.93% 68.32% 45.46% 1.157 45.31% 12
Qwen-VL 54.81% 66.25% 62.91% 60.64% 63.21% 58.67% 83.87% 72.41% 73.90% 64.18% 46.25% 1.192 43.71% 13
LLaVA 46.79% 63.59% 65.30% 64.54% 62.43% 58.61% 80.63% 65.52% 75.83% 64.68% 45.96% 1.125 43.55% 14
TinyGPT-V 21.79% 24.52% 22.13% 28.01% 23.71% 30.36% 29.03% 31.03% 35.40% 32.04% 43.57% 0.701 42.36% 15
InstructBLIP 37.82% 55.36% 55.43% 57.09% 54.29% 49.64% 58.06% 51.72% 61.50% 53.89% 46.54% 1.126 38.96% 16
Otter 35.90% 54.28% 51.65% 51.06% 50.96% 48.42% 70.97% 51.72% 63.21% 53.64% 44.86% 1.027 37.62% 17
IDEFICS-Instruct 37.50% 52.87% 52.84% 51.06% 50.82% 43.93% 64.52% 62.07% 64.06% 50.82% 45.00% 1.180 36.96% 18
MiniGPT-v2 56.73% 56.44% 51.74% 50.00% 54.18% 52.52% 58.06% 44.83% 58.07% 54.36% 31.11% 1.003 35.16% 19
@@ 4.2. Supervised Fine-Tuning
In general, the training of open-source MLLMs (Liu et al., 2023b, a; Zhu et al., 2023) includes two stages: aligning the representation space of the visual backbone and the LLM with million-scale web data (Lu et al., 2022), and visual instruction tuning with a combination of multi-modality datasets (Chen et al., 2023a; Dai et al., 2023). Considering that our purpose is to improve the aesthetic perception ability of the current MLLMs, we directly use the proposed AesMMIT dataset to perform supervised instruction fine-tuning on the models pre-trained on general-purpose visual tasks (Liu et al., 2023a). Following existing works (Wu et al., 2023b; Chen et al., 2023b), supervised instruction fine-tuning not only enables the models to retain their original general knowledge but also facilitates the aesthetics perception capabilities. In this work, to enhance the training efficiency and compare fairly, we freeze the vision model and focus on fine-tuning the projector and the language model. Based on supervised instruction fine-tuning, we implemented three different versions of multi-modality aesthetic expert models.
@@
[…]
5. Experiments
5.1. Implementation Details
Training setting: In this work, we finetune three pre-trained MLLMs based on the constructed AesMMIT dataset in full schedule mode, including LLaVA-1.5-7B (Vicuna-v1.5-7B) (Liu et al., 2023a), LLaVA-1.5-13B (Vicuna-v1.5-13B) (Liu et al., 2023a) and mPLUG-Owl2 (LLaMA-2-7B) (Ye et al., 2023b). In implementation, to ensure fairness, we follow the default hyper-parameters provided by the original models. All models are trained based on 8 NVIDIA Tesla A100 80G GPUs, and the evaluation experiments are conducted on 2 NVIDIA RTX 4090 24G GPUs.
Benchmark: To verify the performance of our proposed AesExpert models on image aesthetics perception, extensive experiments and comparisons are conducted on the AesBench (Huang et al., 2024b), which is a well-designed benchmark for MLLMs on aesthetics perception evaluation. Specifically, AesBench contains 2,800 images and four evaluation criteria designed from four dimensions, including (1) Aesthetic Perception (AesP) focuses on the ability of MLLMs to recognize and understand aesthetic attributes. (2) Aesthetic Empathy (AesE) evaluates the ability of MLLMs to resonate with the emotional aspects conveyed through aesthetic expressions like humans. (3) Aesthetic Assessment (AesA) evaluates the ability of MLLMs to judge aesthetic grades and predict quality scores based on the language description. (4) Aesthetic Interpretation (AesI) involves the ability of MLLMs to interpret and analyze the reasons for aesthetic quality. Each dimension contains 2,800 questions and correct answers. For the first three dimensions, the accuracy of the answers is used to measure the performance of the model, while the AesI is evaluated based on GPT scoring.
@@ 5.2. Performance Comparison
In this section, we compare the performance of our proposed AesExpert models (three versions) with 16 state-of-the-art MLLMs, including the popular GPT-4V (Yang et al., 2023) and Gemini Pro Vision (Google, 2023), as well as 13 state-of-the-art variants with open sources, i.e. LLaVA (LLaMA-2-Chat-7B) (Liu et al., 2023b), LLaVA-1.5-7B (Vicuna-v1.5-7B) (Liu et al., 2023a), LLaVA-1.5-13B (Vicuna-v1.5-13B) (Liu et al., 2023a), ShareGPT4V (Vicuna-v1.5-7B) (Chen et al., 2023a), Q-Instruct (LLaVA-v1.5-7B) (Wu et al., 2023b), mPLUG-Owl2 (LLaMA-2-7B) (Ye et al., 2023b), InstructBLIP (Vicuna-7B) (Dai et al., 2023), MiniGPT-4 (Vicuna-7B) (Zhu et al., 2023), MiniGPT-v2 (LLaMA-2-Chat-7B) (Chen et al., 2023b), IDEFICS-Instruct (LLaMA-7B) (Huggingface, 2023), GLM (ChatGLM-6B) (Du et al., 2022), Otter (MPT-7B) (Li et al., 2023b), TinyGPT-V (Phi-2) (Yuan et al., 2023) and Qwen-VL (QWen-7B) (Bai et al., 2023). Detailed information about these models can be found in (Zhang et al., 2024). The results are listed in Table 3.
From Table 3, we can find that the three versions of AesExpert achieve the top three results. Among them, AesExpert based on LLaVA-1.5-13B achieves the best performance, which is significantly better than the most advanced GPT-4V. For the existing open-source models, ShareGPT4V (Chen et al., 2023a) performs best, but lags behind our AesExpert (L-13B) by more than 10%. These experimental results reveal that the proposed AesExpert models have the best aesthetic perception abilities and highlight the advantage of the constructed AesMMIT dataset for improving multi-modality foundation models.
@@ 5.3. Performance Improvement
In this section, we quantitatively evaluate the aesthetic perception abilities of MLLMs after aesthetic instruction tuning in the four tasks defined by AesBench (Huang et al., 2024b). The experimental results are summarized in Tables 4, 5, 6 and 7, respectively.
@@
Aesthetic Empathy Ability. From Table 5, it is observed that our AesExpert is superior to the baseline models by a large margin (more than 12%), especially on LLaVA-1.5-7B, with performance improvement up to 15.72%. Moreover, the average improvement on Uniqueness is more significant than on other dimensions, indicating that the major concerns for empathy raised by humans in the AesMMIT dataset are related to the uniqueness of perspective. Among the three types of images, artificial intelligence-generated images still obtain the biggest performance improvement (more than 15%). In addition, for the four question types, HOW questions achieve the biggest performance improvement. In summary, these results underscore that AesMMIT can significantly improve the aesthetic empathy abilities of MLLMs, and even our AesExpert based on LLaVA-1.5-7B markedly surpasses the current top-performing GPT-4V (refer to Table 3).
Aesthetic Assessment Ability. The observations from Table 6 underscore that the aesthetic instruction tuning also notably improves the aesthetic assessment ability of MLLMs, especially on the artificial intelligence-generated images with an average improvement of 10.46%. In contrast, the average improvement on artistic image (+1.87) is less significant, implying that, due to the highly abstract nature, the aesthetic assessment of artistic images is still a relatively difficult task. We look forward to better solutions for artistic image in the future.
@@
[…]
Aesthetic Interpretation Ability. Hallucination has been regarded as one of the critical challenges for MLLMs (Han et al., 2024; Huang et al., 2024a), which imagines incorrect details about an image in visual question answering. To alleviate this problem, the proposed AesMMIT dataset is collected from human natural language feedback rather than machine-generated annotations. As can be seen from Table 7, the Precision of interpretation has been significantly improved. For three different baseline MLLMs, significant improvements of 41.50%, 48.10% and 25.90% have been achieved improved, respectively. This result proves that the proposed AesMMIT allows the MLLMs to learn the style of human language and enhance the precision of aesthetic descriptions. In addition, from the overall score, the aesthetic instruction tuning based on AesMMIT significantly improves the aesthetic interpretation ability of MLLMs (average performance improvement of 15.47%), especially for LLaVA-1.5-7B, with a performance improvement up to 19.20%. These results demonstrate that the AesMMIT dataset could significantly benefit the existing MLLMs for obtaining enhanced aesthetic interpretation ability.
@@
5.4. Comparison of Training Data
To verify the effectiveness and necessity of collecting human feedback for improving the aesthetic perception abilities of MLLMs, we further conduct experiments to compare our dataset with the AVA-Comments dataset (Murray et al., 2012), the largest multi-modality dataset in the image aesthetics domain, which contains over 250K images with 1,455K comments. Considering that the AVA-Comments dataset only contains aesthetic descriptions, for fairness, we evaluate the aesthetic interpretation ability across different datasets using the same instruction settings. In addition, we use the AesFeedback subset and the whole AesMMIT dataset to fine-tune the model for comparison, respectively. Figure 5 provides the experimental results, where all experiments adopt the same LLaVA-1.5-7B as the baseline MLLM.

### v3

@@ 3.3. Generating Instruction-following Data
Aesthetic Question Answering: In addition to directly adopting the AesFeedback as aesthetic instruction-following data, we further propose a GPT-assist approach to refine the aesthetic critiques and assemble a larger-scale aesthetic question-answering subset (AesVQA). Inspired by Q-Instruct (Wu et al., 2023b) and ShareGPT (Chen et al., 2023a), we prompt GPT-4 to generate diverse-style open-ended questions and provide corresponding brief answers, where both the questions and answers are based on human aesthetic descriptions. To ensure that MLLMs can handle diverse queries during the interaction, the questions are generated to cover commonly used question types. Specifically, Yes-or-No-style questions are straightforward queries that demand a simple yes or no as an answer, which are mainly used to provide a clear, binary response. What-style questions are leveraged to measure more comprehensive and complex aesthetic perception (e.g. various aesthetic attributes). How-style questions are used to ask more details about aesthetic emotion, aesthetic attributes, and improvement suggestions; Why-style questions are employed to explore the foundational aspects of what makes images aesthetically appealing or unappealing, revealing the rationale of beauty and visual appeal. To ensure the diversity of instructions, following the existing works (Liu et al., 2023a; Lu et al., 2022), besides the direct answers, we also generate several wrong alternatives for the questions and convert them into an additional multi-choice question format (Wu et al., 2024b). Through the above process, we obtain the final AesMMIT dataset, which consists of 409K multi-typed instructions, aiming to activate stronger aesthetic capabilities. Figure 3 shows the frequently occurring words in the proposed AesMMIT dataset. It can be observed that most nouns, adjectives and adverbs are related to image aesthetic description, which is quite different from common semantic-based tasks.
@@ 4.1. Model Architecture
The proposed AesExpert model follows the design of LLaVA-1.5 (Liu et al., 2023a), which includes three components: (1) A visual model based on the CLIP-ViT-L14 (Radford et al., 2021) with an input size of $\times×336$, which converts the input image into 576 tokens. (2) A visual-language projector based on two-layer multi-layer perception (MLP), which is employed to connect the visual modality and language modality. (3) A language model based on the open-sourced Vicuna-v1.5 (Zheng et al., 2024), which is used to interpret instructions and generate responses. In this work, we build the AesExpert models based on two variants of LLaVA-1.5, including LLaVA-1.5-7B and LLaVA-1.5-13B. In addition, we also introduce the mPLUG-Owl2 (Ye et al., 2023b) as the backbone to validate the aesthetic perception performance based on the AesMMIT dataset. The structures for these multi-modality foundation models are summarized in Table 2.
@@
mPLUG-Owl2 60.90% 70.57% 68.30% 62.77% 67.89% 65.60% 77.42% 65.52% 78.07% 69.89% 50.57% 1.182 76.64% 8
Q-Instruct 66.03% 74.48% 73.68% 68.09% 72.61% 68.64% 83.86% 75.86% 80.00% 72.68% 52.86% 1.020 75.04% 9
LLaVA-1.5-7B 53.85% 70.16% 67.40% 59.93% 66.32% 62.49% 80.65% 75.85% 78.93% 68.32% 45.46% 1.157 73.95% 10
Qwen-VL 54.81% 66.25% 62.91% 60.64% 63.21% 58.67% 83.87% 72.41% 73.90% 64.18% 46.25% 1.192 73.21% 11
LLaVA 46.79% 63.59% 65.30% 64.54% 62.43% 58.61% 80.63% 65.52% 75.83% 64.68% 45.96% 1.125 71.39% 12
InstructBLIP 37.82% 55.36% 55.43% 57.09% 54.29% 49.64% 58.06% 51.72% 61.50% 53.89% 46.54% 1.126 66.83% 13
IDEFICS-Instruct 37.50% 52.87% 52.84% 51.06% 50.82% 43.93% 64.52% 62.07% 64.06% 50.82% 45.00% 1.180 66.16% 14
Otter 35.90% 54.28% 51.65% 51.06% 50.96% 48.42% 70.97% 51.72% 63.21% 53.64% 44.86% 1.027 63.04% 15
MiniGPT-v2 56.73% 56.44% 51.74% 50.00% 54.18% 52.52% 58.06% 44.83% 58.07% 54.36% 31.11% 1.003 59.99% 16
GLM 55.77% 54.61% 51.25% 48.94% 52.96% 53.13% 70.97% 44.83% 55.29% 53.96% 37.79% 0.932 59.48% 17
MiniGPT-4 39.42% 41.31% 42.67% 44.33% 41.93% 39.78% 38.71% 24.14% 39.04% 39.35% 38.57% 0.999 54.94% 18
TinyGPT-V 21.79% 24.52% 22.13% 28.01% 23.71% 30.36% 29.03% 31.03% 35.40% 32.04% 43.57% 0.701 42.36% 19
@@ 4.2. Supervised Fine-Tuning
In general, the training of open-sourced MLLMs (Liu et al., 2023b, a; Zhu et al., 2023a) includes two stages: aligning the representation space of the visual backbone and the LLM with million-scale web data (Lu et al., 2022), and visual instruction tuning with a combination of multi-modality datasets (Chen et al., 2023a; Dai et al., 2023). Considering that our purpose is to improve the aesthetic perception ability of the current MLLMs, we directly use the proposed AesMMIT dataset to perform supervised instruction fine-tuning on the models pre-trained on general-purpose visual tasks (Liu et al., 2023a). Following existing works (Wu et al., 2023b; Chen et al., 2023b), supervised instruction fine-tuning not only enables the models to retain their original general knowledge but also facilitates the aesthetics perception capabilities. In this work, to enhance the training efficiency and compare fairly, we freeze the vision model and focus on fine-tuning the projector and the language model. Based on supervised instruction fine-tuning, we implemented three different versions of multi-modality aesthetic expert models.
5. Experiments
5.1. Implementation Details
Training setting: In this work, we finetune three pre-trained MLLMs based on the constructed AesMMIT dataset in full schedule mode, including LLaVA-1.5-7B (Vicuna-v1.5-7B) (Liu et al., 2023a), LLaVA-1.5-13B (Vicuna-v1.5-13B) (Liu et al., 2023a) and mPLUG-Owl2 (LLaMA-2-7B) (Ye et al., 2023b). In implementation, to ensure fairness, we follow the default hyper-parameters provided by the original models. All models are trained based on the GPU server with 8 NVIDIA Tesla A100 80G, and the evaluation experiments are conducted on the GPU server with 2 NVIDIA RTX 4090 24G.
Benchmark: To verify the performance of our proposed AesExpert models on image aesthetics perception, extensive experiments and comparisons are conducted on the AesBench (Huang et al., 2024c), which is a well-designed benchmark for MLLMs on aesthetics perception evaluation. Specifically, AesBench contains 2,800 images and four evaluation criteria designed from four dimensions, including (1) Aesthetic Perception (AesP) focuses on the ability of MLLMs to recognize and understand aesthetic attributes. (2) Aesthetic Empathy (AesE) evaluates the ability of MLLMs to resonate with the emotional aspects conveyed through aesthetic expressions like humans. (3) Aesthetic Assessment (AesA) evaluates the ability of MLLMs to judge image aesthetics quality. (4) Aesthetic Interpretation (AesI) involves the ability of MLLMs to interpret and analyze the reasons for aesthetic quality. Each dimension contains 2,800 questions and correct answers. For the first three dimensions, the accuracy of the answers is used to measure the performance of the model, while the AesI is evaluated based on GPT scoring.
@@ 5.2. Performance Comparison
In this section, we compare the performance of our proposed AesExpert models (three versions) with 16 state-of-the-art MLLMs, including the popular GPT-4V (Yang et al., 2023) and Gemini Pro Vision (Google, 2023), as well as 14 state-of-the-art variants with open sources, i.e. LLaVA (LLaMA-2-Chat-7B) (Liu et al., 2023b), LLaVA-1.5-7B (Vicuna-v1.5-7B) (Liu et al., 2023a), LLaVA-1.5-13B (Vicuna-v1.5-13B) (Liu et al., 2023a), ShareGPT4V (Vicuna-v1.5-7B) (Chen et al., 2023a), Q-Instruct (LLaVA-v1.5-7B) (Wu et al., 2023b), mPLUG-Owl2 (LLaMA-2-7B) (Ye et al., 2023b), InstructBLIP (Vicuna-7B) (Dai et al., 2023), IDEFICS-Instruct (LLaMA-7B) (Huggingface, 2023), Otter (MPT-7B) (Li et al., 2023b), Qwen-VL (QWen-7B) (Bai et al., 2023), GLM (ChatGLM-6B) (Du et al., 2022), MiniGPT-v2 (LLaMA-2-Chat-7B) (Chen et al., 2023b), TinyGPT-V (Phi-2) (Yuan et al., 2023) and MiniGPT-4 (Vicuna-7B) (Zhu et al., 2023a). Detailed information about these models can be found in (Zhang et al., 2024). The experimental results are listed in Table 3.
From Table 3, we can find that the three versions of AesExpert achieve the top three results. Among them, AesExpert based on LLaVA-1.5-13B achieves the best performance, which is significantly better than the most advanced GPT-4V. For the existing open-sourced models, ShareGPT4V (Chen et al., 2023a) performs best, but lags behind our AesExpert (L-13B) by more than 10% on overall score. These experimental results reveal that the proposed AesExpert models have the best aesthetic perception abilities and highlight the advantage of the constructed AesMMIT dataset for improving multi-modality foundation models.
@@ 5.3. Performance Improvement
In this section, we quantitatively evaluate the aesthetic perception abilities of MLLMs after aesthetic instruction tuning in the four tasks defined by AesBench (Huang et al., 2024c). The experimental results are summarized in Tables 4, 5, 6 and 7, respectively.
@@
Aesthetic Empathy Ability. From Table 5, it is observed that our AesExpert is superior to the baseline models by a large margin (more than 12%), especially on LLaVA-1.5-7B, with performance improvement up to 15.72%. Among the four dimensions, Uniqueness usually has significant improvement, indicating that the major concerns for empathy raised by humans in the AesMMIT dataset are related to the uniqueness of perspective. Among the three types of images, artificial intelligence-generated images still obtain the biggest performance improvement (more than 15%). In addition, for the four question types, HOW questions achieve the biggest performance improvement. In summary, these results underscore that AesMMIT can significantly improve the aesthetic empathy abilities of MLLMs, and even our AesExpert based on the LLaVA-1.5-7B model markedly surpasses the current top-performing GPT-4V (refer to Table 3).
Aesthetic Assessment Ability. The observations from Table 6 underscore that the aesthetic instruction tuning also notably improves the aesthetic assessment ability of MLLMs, especially on the artificial intelligence-generated images. In contrast, the average improvement on artistic image is less significant, implying that, due to the highly abstract nature, the aesthetic assessment of artistic images is still a relatively difficult task. We look forward to better solutions for artistic image in the future.
@@
Aesthetic Interpretation Ability. Hallucination has been regarded as one of the critical challenges for MLLMs (Han et al., 2024; Huang et al., 2024b), which imagines incorrect details about an image in visual question answering. To alleviate this problem, the proposed AesMMIT dataset is collected from human natural language feedback rather than machine-generated annotations. As can be seen from Table 7, the Precision of interpretation has been significantly improved. For three different baseline MLLMs, significant improvements of 41.50%, 44.60% and 25.90% have been achieved improved, respectively. This result proves that the proposed AesMMIT allows the MLLMs to learn the style of human language and enhance the precision of aesthetic descriptions. In addition, from the overall score, the aesthetic instruction tuning based on AesMMIT significantly improves the aesthetic interpretation ability of MLLMs, especially for LLaVA-1.5-7B and mPLUG-Owl2 with performance improvement of 16.00% and 15.40%. These results demonstrate that the AesMMIT dataset could significantly benefit the existing MLLMs for obtaining enhanced aesthetic interpretation ability.
5.4. Comparison of Training Data
To verify the effectiveness and necessity of collecting human feedback for improving the aesthetic perception abilities of MLLMs, we further conduct experiments to compare our dataset with the AVA-Comments dataset (Murray et al., 2012), the largest multi-modality dataset in image aesthetics assessment, which contains over 250K images with 1,455K comments. Considering that the AVA-Comments dataset only contains aesthetic descriptions, for fairness, we evaluate the aesthetic interpretation ability across different datasets using the same instruction settings. In addition, we use the AesFeedback subset and the whole AesMMIT dataset to fine-tune the model for comparison, respectively. Figure 4 provides the experimental results, where all experiments adopt the same LLaVA-1.5-7B as the baseline MLLM.